OpenStudy (anonymous):

find the distance between the two points. round to the nearest tenth if necessary (2,2), (8,10) A. 6 B. 10 C. 8 D. 100

OpenStudy (anonymous):

@freckles

OpenStudy (freckles):

\[\text{ The distance between the points } (a,b) \text{ and } (c,d) \\ \text{ can be found by computing } \\ \sqrt{(a-c)^2+(b-d)^2} \]

OpenStudy (freckles):

so first let's do this with little bitty steps the difference of your x's is 2-8 which is? the difference of your y's is 2-10 which is?

OpenStudy (anonymous):

-6 and -8

OpenStudy (freckles):

great now once we find the differences we square each

OpenStudy (freckles):

(-6)^2=? (-8)^2=?

OpenStudy (anonymous):

um.....

OpenStudy (freckles):

(-6)^2=(-6)(-6)=? (-8)^2=(-8)(-8)=?

OpenStudy (freckles):

\[\text{ The distance between the points } (a,b) \text{ and } (c,d) \\ \text{ can be found by computing } \\ \sqrt{(a-c)^2+(b-d)^2}\] \[\text{ The distance between the points } (2,2) \text{ and } (8,10) \\ \text{ can be found by computing } \\ \sqrt{(2-8)^2+(2-10)^2}\] \[\sqrt{(2-8)^2+(2-10)^2} \\ \sqrt{(-6)^2+(-8)^2} \\ \sqrt{36+64}\] can you finish from here?

OpenStudy (anonymous):

it is b

OpenStudy (freckles):

yep the sqrt(100) is 10